# Coherent Optical Modems for Full-Wavefield Lidar

### v1

@@ Abstract.
The advent of the digital age has driven the development of coherent optical modems—devices that modulate the amplitude and phase of light in multiple polarization states. These modems transmit data through fiber optic cables that are thousands of kilometers in length at data rates exceeding one terabit per second. This remarkable technology is made possible through near-THz-rate programmable control and sensing of the full optical wavefield. While coherent optical modems form the backbone of telecommunications networks around the world, their extraordinary capabilities also provide unique opportunities for imaging. Here, we introduce full-wavefield lidar: a new imaging modality that repurposes off-the-shelf coherent optical modems to simultaneously measure distance, axial velocity, and polarization. We demonstrate this modality by combining a 74 GHz-bandwidth coherent optical modem with free-space coupling optics and scanning mirrors. We develop a time-resolved image formation model for this system and formulate a maximum-likelihood reconstruction algorithm to recover depth, velocity, and polarization information at each scene point from the modem’s raw transmitted and received symbols. Compared to existing lidars, full-wavefield lidar promises improved mm-scale ranging accuracy from brief, microsecond exposure times, reliable velocimetry, and robustness to intererence from ambient light or other lidar signals.
@@ 1. Introduction
The extreme abilities of these devices to manipulate and sense light within a fiber raise the question: how can we leverage off-the-shelf optical modems to advance the state of the art in free-space imaging? As a first step toward addressing this question, we introduce full-wavefield lidar (FWL), a new lidar sensing modality for simultaneous measurement of distance, axial velocity, and two orthogonal linear polarization states using coherent optical modems.
@@ 2.1. Incoherent Lidar
Most commercial lidar systems operate on a principle of incoherent detection. These lidars modulate the intensity of light to recover scene geometry by measuring phase delays of a sinusoidally modulated signal (Coddington et al., 2009; Schuhler et al., 2006) or propagation delays of emitted and backscattered laser pulses (Rapp et al., 2020; Takeuchi et al., 1983; Lin and Liu, 2004). Incoherent lidars can also capture polarization information of backscattered light (Baek and Heide, 2021, 2022). However, incoherent detection schemes are sensitive to interference from other light sources or ambient light (e.g., from other lidars or the sun). While incoherent continuous wave systems can measure velocity from phase shifts due to the Doppler effect (Hu et al., 2022; Heide et al., 2015), pulsed systems are not sensitive to phase information and cannot be used for velocimetry in the same fashion. FWL recovers accurate depth with 1 $\mu$s exposure times that are 10,000$\times$ shorter than that of incoherent, intensity-modulated depth sensors (e.g., the Kinect Azure (Bamji et al., 2018), which uses $\approx$10 ms exposure times—see Figure 2).
@@ 2.2. Coherent Lidar
Perhaps the most common type of coherent lidar is based on a frequency-modulated continuous wave (FMCW) transmit signal (Zheng, 2005). Specifically, FMCW lidars transmit a “chirp” signal whose optical frequency increases linearly with time. The depth resolution is tied to the frequency tuning range of the laser source, and the maximum range depends on the coherence length of the laser (Amann, 1992). Still, FMCW lidars are sensitive to interference from other frequency-modulated lidars and require transmitting multiple chirps to unambiguously resolve range and velocity.
@@
Many of the drawbacks associated with FMCW lidar are mitigated with random modulation continuous wave (RMCW) lidar (Takeuchi et al., 1983). This modality uses a random phase and amplitude encoding that suppresses interference (Hwang and Lee, 2020), and enables unambiguous ranging and velocimetry. However, there are considerable implementation challenges. The range resolution is determined, in part, by the modulation speed and sample rate—typically tens of GHz to achieve mm-scale resolution. Thus, few examples of this modality exist in the literature due to the significant hardware requirements related to ultrafast sample rates and the computational challenge of modeling Doppler shifts, laser phase noise, speckle, and polarization changes induced by scattering.
@@ 3.1. Coherent Modulation and Demodulation
(1) $\mathbf{{X}}(t)=\sum_{n=0}^{N-1}\mathbf{{X}}_{n}\>\mathsf{rect}\bigg{(}t-nT\bigg{)}*\mathsf{{B}}(t),$
@@
We note that this section provides a simplified description of the coherent modulation and demodulation procedure. In practice, optical modems contain additional optical and digital processing stages to ensure that the local oscillator is locked to the transmit laser and that the received signal is sampled with the correct timing. Further, we note that the Gaussian noise model neglects secondary effects such as optical fiber non-linearities (Singh and Singh, 2007). For a detailed treatment of coherent detection in optical modems, please refer to Ip et al. (2008).
@@ 3.2. Repurposing Coherent Optical Modems for Imaging
The received and demodulated electric field $\mathbf{{E}}_{\text{RX}}(t)$ is corrupted by three main effects resulting from propagation to the scene and back. First, the demodulated electric field is time-shifted relative to the transmit signal due to the propagation delay to the target and back. Second, if illuminating a moving target, the field is frequency-shifted due to the Doppler effect (McManamon, 2015). Finally, the field is distorted by attenuation and changes in the polarization state due to the surface reflection. The received electric field can therefore be modeled as
@@
The terms $\tau$ and $\nu$ denote the propagation delay and frequency shifts of the demodulated electric field. The Jones matrix $\mathbf{{R}}$ is a $2\times 2$ complex-valued matrix that describes how the transmitted electric field is transformed by a polarization-dependent attenuation and rotation induced by the properties of the target surface and propagation through the fiber (Gordon and Kogelnik, 2000)).
@@ 4. Joint Estimation of Depth, Velocity, and Polarization
where $\Delta_{\text{min}}$ is the minimum delay due to free-space propagation, ignoring internal reflections in the optical modem [^2] We assume that the values $\Delta$ associated with internal reflections are calibrated a priori..
@@ 4.1. Implementation Details
Table 1. Evaluation of depth precision. We scan a planar surface and measure deviation of the measurements to a plane fitted to the surface. We compare to performance using the generalized matched filter; note that we omit the TV regularizer for this evaluation to assess per-pixel precision.
mean depth error (mm)\,$\downarrow$ % of pixels with depth error ¡ 2 mm\,$\uparrow$ % of pixels with depth error ¡ 6 mm\,$\uparrow$
method joint estimation gen. matched filter joint estimation gen. matched filter joint estimation gen. matched filter
FWL 4.43 9.93 65.20 64.77 98.50 97.78
dual-polarization phase 9.31 24.99 57.16 55.40 97.91 94.92
dual-polarization and amplitude 19.08 39.51 58.59 55.72 95.96 91.92
single-polarization phase and amplitude 30.19 46.95 47.85 45.83 93.75 90.46
@@
A photo and illustration of the hardware prototype are shown in Figure 8. We use a Ciena WaveLogic 5n modem with a sampling frequency of $741/T=$ 74 GHz (i.e., an optical path resolution of $4$ mm or depth resolution of $2$ mm); the laser operates at a wavelength of 1550 nm. We communicate with the modem over its QSFP-DD electrical interface (Ghiasi, 2023) to program the transmit sequence and read out the measured data. The transmitted sequence length, limited by finite modem memory, is set to $\approx 2^{16}$ symbols, providing a maximum unambiguous range of approximately 130 meters. We use an Erbium-doped fiber amplifier (EDFA (Becker et al., 1999)) to boost the power of the emitted laser light from 1 mW up to a maximum of 100 mW; all experiments use an eye-safe transmit laser power of 2 mW unless otherwise specified. Another EDFA is used as a pre-amplifier to boost the power of the received light up to the level expected by the modem (roughly 1 mW).
@@ 5.2. Imaging Demonstrations
In Figure 1 we show a scene illuminated by sunlight (through a window) containing static objects and a spinning hemisphere. The radial velocity of the hemisphere is estimated to be a maximum of 25 m/s, which agrees with estimates produced using a high speed camera (see supplement). Depth is recovered despite the presence of strong ambient light from the sun, which illuminates the scene through the window (see Figure 11 for a point cloud reconstruction).
@@
In Figure 11 (a), we demonstrate reconstructing objects with subsurface scattering and detailed geometry and provide a comparison to generalized matched filtering using optical powers of 2 mW and 80 mW. In each case the qualitative results in Figure 11 show robust reconstruction of the scene geometry. We see similar trends for a larger room-sized scene captured with FWL.
@@
In Figure 11 (c), we show an example scene where a statuette is placed behind a translucent surface. Our optimization procedure recovers both surfaces as shown in the reconstructed point cloud.
@@ 6. Discussion
Coherent optical modems are a promising solution to make coherent lidar more accessible to researchers and practitioners. Still, some barriers remain to widespread adoption of this technology. For example, using fast optical modems (operating at tens to hundreds of GHz) requires domain knowledge to program and read out the transmitted and received waveforms. Currently, our hardware interface to the modem requires around one second to transfer data to a computer after each exposure; this limits the acquisition speeds of the current system to more than a second per scan point and thus limits the overall scan resolution. In future work we hope to address these obstacles and to demonstrate real-time capture of depth and velocity. FWL may also offer benefits for other imaging scenarios, such as in the presence of scattering media, where sensitivity to motion, depth, and polarization could help to isolate unscattered light.
@@ Acknowledgments
We thank Ciena Corporation for loaning the coherent optical modem and Praneeth Chakravarthula for helpful discussions. DBL and KNK acknowledge support from Ciena Corporation, and NSERC under the RGPIN, RTI, and Alliance programs. DBL also acknowledges support from the Canada Foundation for Innovation and the Ontario Research Fund. PM was supported by Mitacs through the Mitacs Accelerate program.

### v2

@@ Abstract.
The advent of the digital age has driven the development of coherent optical modems—devices that modulate the amplitude and phase of light in multiple polarization states. These modems transmit data through fiber optic cables that are thousands of kilometers in length at data rates exceeding one terabit per second. This remarkable technology is made possible through near-THz-rate programmable control and sensing of the full optical wavefield. While coherent optical modems form the backbone of telecommunications networks around the world, their extraordinary capabilities also provide unique opportunities for imaging. Here, we repurpose off-the-shelf coherent optical modems to introduce full-wavefield lidar: a type of random modulation continuous wave lidar that simultaneously measures depth, axial velocity, and polarization. We demonstrate this modality by combining a 74 GHz-bandwidth coherent optical modem with free-space coupling optics and scanning mirrors. We develop a time-resolved image formation model for this system and formulate a maximum-likelihood reconstruction algorithm to recover depth, velocity, and polarization information at each scene point from the modem’s raw transmitted and received symbols. Compared to existing lidars, full-wavefield lidar promises improved mm-scale ranging accuracy from brief, microsecond exposure times, reliable velocimetry, and robustness to interference from ambient light or other lidar signals.
coherent lidar, coherent optical modems, 3D imaging, Doppler imaging, polarization imaging
submissionid: 856journalyear: 2024copyright: acmlicensedconference: SIGGRAPH Asia 2024 Conference Papers; December 3–6, 2024; Tokyo, Japanbooktitle: SIGGRAPH Asia 2024 Conference Papers (SA Conference Papers ’24), December 3–6, 2024, Tokyo, Japandoi: 10.1145/3680528.3687663isbn: 979-8-4007-1131-2/24/12ccs: Computing methodologies Computer visionccs: Computing methodologies Computational photographyccs: Computing methodologies 3D imaging
@@ 1. Introduction
The extreme abilities of these devices to manipulate and sense light within a fiber raise the question: how can we leverage off-the-shelf optical modems to advance the state of the art in free-space imaging? As a first step toward addressing this question, we introduce full-wavefield lidar (FWL). This sensing modality falls broadly within the class of random modulation continuous wave (RMCW) lidar; however, compared to conventional RMCW techniques (Sambridge et al., 2021) that use binary phase (Bashkansky et al., 2004) or amplitude (Ai et al., 2011) modulation, our approach employs controllable modulation of the full coherent wavefield across two polarization states. Further, our FWL reconstruction algorithm provides simultaneous estimates of depth and velocity without requiring specialized encoding schemes or trading off bandwidth (Banzhaf and Waldschmidt, 2021).
@@ 2.1. Incoherent Lidar
Most commercial lidar systems operate on a principle of incoherent detection. These lidars modulate the intensity of light to recover scene geometry by measuring phase delays of a sinusoidally modulated signal (Coddington et al., 2009; Schuhler et al., 2006) or propagation delays of emitted and backscattered laser pulses (Rapp et al., 2020; Takeuchi et al., 1983; Lin and Liu, 2004). Incoherent lidars can also capture polarization information of backscattered light (Baek and Heide, 2021, 2022). However, incoherent detection schemes are sensitive to interference from other light sources or ambient light (e.g., from other lidars or the sun). While incoherent continuous wave systems can measure velocity from phase shifts due to the Doppler effect (Hu et al., 2022; Heide et al., 2015), pulsed systems are not sensitive to phase information and cannot be used for velocimetry in the same fashion. FWL recovers accurate depth with 1 $\mu$s exposure times that are 10,000$\times$ shorter than that of incoherent, intensity-modulated depth sensors (e.g., the Kinect Azure (Bamji et al., 2018), which uses $\approx$10 ms exposure times—see Figure 2). Also, similar to pulsed systems (O’Toole et al., 2017) and some continuous-wave systems (Kadambi et al., 2013; Peters et al., 2015), our approach recovers time-resolved profiles of backscattered light.
@@ 2.2. Coherent Lidar
Perhaps the most common type of coherent lidar is based on a frequency-modulated continuous wave (FMCW) transmit signal (Zheng, 2005). Specifically, FMCW lidars transmit a “chirp” signal whose optical frequency increases linearly with time. While one advantage of FMCW lidar is the simplicity of its modulation scheme (i.e., the linear frequency sweep), the system performance is strongly coupled to the modulation waveform. FMCW depth resolution is tied to the bandwidth of the linear frequency sweep, and the maximum depth and velocity resolution depend on the duration of the sweep (Behroozpour et al., 2017). For many swept-source lasers, there is limited flexibility to adjust these parameters (e.g., beyond certain preset configurations). In contrast, our system’s performance is easily configurable and largely decoupled from the modulation waveform: the depth resolution is tied to the sample rate (74 GHz for our system), and the maximum depth and the velocity resolution depend on the exposure time—each of which can be continuously controlled. Our setup is also more general as it measures the response of a scene to arbitrary, programmable wavefields. Additionally, FMCW lidars are sensitive to interference from other frequency-modulated lidars (Hwang et al., 2020). RMCW techniques, including our proposed FWL system, are less susceptible to interference than FMCW because the transmit waveforms are likely to be orthogonal to any interfering signal (Chen et al., 2023; Hwang and Lee, 2020).
@@
FWL can be considered as a type of RMCW lidar (Takeuchi et al., 1983); however, different from conventional RMCW lidar, we do not use a binary modulation of phase (Bashkansky et al., 2004; Sambridge et al., 2021) or amplitude (Ai et al., 2011; Takeuchi et al., 1986). Instead, we use discrete, pseudorandom modulation of both amplitude and phase in two polarization states. In this sense, our work is close to “true random” or “chaos” lidars that use amplified spontaneous emission noise (Hwang and Lee, 2020) or chaotic microcombs (Chen et al., 2023) to induce random fluctuations in the amplitude of emitted light. However, unlike FWL, their output waveforms cannot be programmably controlled since they are governed by stochastic phenomena.
Many of the drawbacks associated with FMCW lidar (e.g., related to configurability or interference as discussed above) can be mitigated with RMCW lidar—though there are considerable implementation challenges. The range resolution is determined, in part, by the modulation speed and sample rate—typically tens of GHz to achieve mm-scale resolution. Thus, few examples of this modality exist in the literature due to the significant hardware requirements related to ultrafast sample rates and the computational challenge of modeling Doppler shifts, laser phase noise, speckle, and polarization changes induced by scattering.
@@ 3.1. Coherent Modulation and Demodulation
(1) $\vspace*{-2pt}\mathbf{{X}}(t)=\sum_{n=0}^{N-1}\mathbf{{X}}_{n}\>\mathsf{rect}\bigg{(}t-nT\bigg{)}*\mathsf{{B}}(t),$
@@
We note that this section provides a simplified description of the coherent modulation and demodulation procedure. In practice, optical modems contain additional optical and digital processing stages to ensure that the local oscillator is locked to the transmit laser and that the received signal is sampled with the correct timing. While shot noise generally follows a Poisson distribution, here it is approximated by a Gaussian distribution. Further, we note that the Gaussian noise model neglects secondary effects such as optical fiber non-linearities (Singh and Singh, 2007). For a detailed treatment of coherent detection in optical modems, please refer to Ip et al. (2008).
@@ 3.2. Repurposing Coherent Optical Modems for Imaging
The received and demodulated electric field $\mathbf{{E}}_{\text{RX}}(t)$ is altered by three main effects resulting from propagation to the scene and back. First, the demodulated electric field is time-shifted relative to the transmit signal due to the propagation delay to the target and back. Second, if illuminating a moving target, the field is frequency-shifted due to the Doppler effect (McManamon, 2015). Finally, the field is distorted by attenuation and changes in the polarization state due to the surface reflection. The received electric field can therefore be modeled as
@@
The terms $\tau$ and $\nu$ denote the propagation delay and frequency shifts of the demodulated electric field. The Jones matrix $\mathbf{{R}}$ is a $2\times 2$ complex-valued matrix that describes how the transmitted electric field is transformed by a polarization-dependent attenuation and rotation induced by the properties of the target surface and propagation through the fiber (Gordon and Kogelnik, 2000)). We will later show that the optical modem’s measurement of two polarization channels enables recovery of the Jones matrix.
@@ 4. Joint Estimation of Depth, Velocity, and Polarization
where $\Delta_{\text{min}}$ is the minimum delay due to free-space propagation, ignoring internal reflections in the optical modem. [^2] We assume that the values $\Delta$ associated with internal reflections are calibrated a priori.
@@ 4.1. Implementation Details
A photo and illustration of the hardware prototype are shown in Figure 8. We use a Ciena WaveLogic 5n modem with a sampling frequency of $741/T=$ 74 GHz (i.e., an optical path resolution of $4$ mm or depth resolution of $2$ mm); the laser operates at a wavelength of 1550 nm. We communicate with the modem over its QSFP-DD electrical interface (Ghiasi, 2023) to program the transmit sequence and read out the measured data. The transmitted sequence length, limited by finite modem memory, is set to $\approx 2^{16}$ symbols, providing a maximum unambiguous range of approximately 130 meters. We use an Erbium-doped fiber amplifier (EDFA (Becker et al., 1999)) to boost the power of the emitted laser light from 1 mW up to a maximum of 100 mW; all experiments use an eye-safe transmit laser power of 2 mW unless otherwise specified. Another EDFA is used as a pre-amplifier to boost the power of the received light up to the level expected by the modem ($\approx$1 mW).
@@ 5.2. Imaging Demonstrations
In Figure 1 we show a scene illuminated by sunlight (through a window) containing static objects and a spinning hemisphere. The radial velocity of the hemisphere is estimated to be a maximum of 25 m/s, which agrees with estimates produced using a high speed camera (see supplement). Depth is recovered despite the presence of strong ambient light from the sun, which illuminates the scene through the window (see Figure 11 (e)).
@@
In Figure 11 (a), we demonstrate reconstructing objects with subsurface scattering and detailed geometry and provide a comparison to generalized matched filtering using optical powers of 2 mW and 80 mW. In each case the qualitative results in Figure 11 show robust reconstruction of the scene geometry. We see similar trends for a larger room-sized scene.
@@
In Figure 11 (c), we show an example scene where a statuette is placed behind a translucent surface. We recover both surfaces by estimating the time delays of the two Jones matrices with the greatest Frobenius norms.
@@ 6. Discussion
Coherent optical modems are a promising solution to make coherent lidar more accessible to researchers and practitioners. Still, some barriers remain to widespread adoption of this technology. For example, using fast optical modems (operating at tens to hundreds of GHz) requires domain knowledge to program and read out the transmitted and received waveforms. Currently, our hardware interface to the modem requires $\sim$1 second to transfer data to a computer after each exposure; this limits the acquisition speeds of the current system to more than a second per scan point and thus limits the overall scan resolution. In future work we hope to address these obstacles and to demonstrate real-time capture of depth and velocity. FWL may also offer benefits for other imaging scenarios, such as in the presence of scattering media, where sensitivity to motion, depth, and polarization could help to isolate unscattered light.
@@ Acknowledgments
We thank Ciena Corporation for loaning the optical modem and Praneeth Chakravarthula for helpful discussions. DBL and KNK acknowledge support from Ciena Corporation, and NSERC under the RGPIN, RTI, and Alliance programs. DBL also acknowledges support from the Canada Foundation for Innovation and the Ontario Research Fund. PM was supported by the Mitacs Accelerate program. MW was supported by an XSeed grant and TRANSFORM HF.